40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 24 15:24:13 2021, gautam, Update, LSC, Notes on tests 
    Reply  Wed Mar 24 22:54:49 2021, gautam, Update, LSC, New day, new problems PRMI3f_noArmssensMat.pdfREFL55_whtGainStepping.pngREFL55_whtGainStepping2.png
       Reply  Mon Mar 29 19:32:46 2021, gautam, Update, LSC, REFL55 whitening checkout REFL55wht.png
Message ID: 15958     Entry time: Wed Mar 24 15:24:13 2021     Reply to this: 15960
Author: gautam 
Type: Update 
Category: LSC 
Subject: Notes on tests 

For my note-taking:

  1. Lock PRMI with ITMs as the MICH actuator. Confirm that the MICH-->PRCL contribution cannot be nulled. ✅  [15960]
  2. Lock PRMI on REFL165 I/Q. Check if transition can be made smoothly to (and from?) REFL55 I/Q.
  3. Lock PRMI. Turn sensing lines on. Change alignment of PRM / BS and see if we can change the orthogonality of the sensing.
  4. Lock PRMI. Put a razor blade in front of an out-of-loop photodiode, e.g. REFL11 or REFL33. Try a few different masks (vertical half / horizontal half and L/R permutations) and see if the orthogonality (or lack thereof) is mask-dependent.
  5. Double check the resistance/inductance of the PRM OSEMs by measuring at 1X4 instead of flange. ✅  [15966]
  6. Check MC spot centering.

If I missed any of the tests we discussed, please add them here.

ELOG V3.1.3-